OpenStudy (liciagrl911):

Will give MEDALS! May someone please explain distributive properties to me?

OpenStudy (mathstudent55):

The general rules are: \(a(b + c) = ab + ac\) Distributive property of multiplication over addition \(a(b - c) = ab - ac\) Distributive property of multiplication over subtraction

OpenStudy (mathstudent55):

Here are some examples with numbers and variables to show you how it works. \(5(3 + 4) = 5 \times 3 + 5 \times 4 = 15 + 20 = 35\) \(2(5 - 3) = 2 \times 5 - 2 \times 3 = 10 - 6 = 4\) \(4(x + 5) = 4x + 4 \times 5 = 4x + 20\) \(2x(4x - y) = (2x)(4x) - (2x)(y) = 8x^2 - 2xy\)

OpenStudy (liciagrl911):

How exactly do I do it? Do you think you could break it down for me?

OpenStudy (mathstudent55):

Do you have a specific problem or a question with an example?

OpenStudy (mathstudent55):

The idea is multiply the number outside the parentheses by the two numbers inside the parentheses.